OpenStudy (anonymous):

Find the derivative of f(x)=1/sqrt(x) using the DEFINITION OF THE DERIVATIVE. (please show ALL steps)

OpenStudy (lacypennelll):

t(x+h) = 1/√(x+h); [f(x+h) - f(x)]/h = [1/√(x+h)] -1/√x]/h = [√x - √(x+h)]/[h√x(x+h)] Now multiply numerator and denominator by √x + √(x+h) Let me work them separately as the algebra gets messy: Numerator: [√x - √(x+h)][√x + √(x+h)] = x - x - h = -h and this is why we multiply by √x + √(x+h): the -h in the numerator will cancel the h in the denominator Denominator: [h√x(x+h)][√x + √(x+h)] = h[x√(x+h) +(x+h)√x] Numerator/Denominator: -h/{h[x√(x+h) +(x+h)√x] } = -1/[x√(x+h) +(x+h)√x] Let h ->0 and get -1/(2x√x) = -1/(2x^3/2)

OpenStudy (anonymous):

thanks